A sailboat costs $30,357. You pay 10% down and amortize the rest with equal monthly payments over a 8-year period. If you must pay 7.8% compounded monthly, what is your monthly payment? How much interest will you pay?

Monthly payments: $____

Interest: $____

(Please square, label, or circle the answer)
Answer #1

a)

Amount financed = Cost (1 - % down payment]

                          = 30357 [1 - .10]

                         = 30357 *.9

                         = 27321.3 (rounded to 27321]

Monthly rate = 7.8/12= .65%

Number of months over 8 years = 8 *12 months a year = 96

Monthly payment =Amount financed /PVA.65%,96

                           = 27321 / 71.25049

                           = 383.45

#Find present value annuity factor using financial calculator where i=.65% ,n= 96 ,PMT =1 and then press PV

b)Total Payment made = 383.45 *96 = 36811.2   (rounded to 36811)

Interest = total payment -Amount financed

               = 36811 - 27321

                = 9490